# Data-Efficient Quantum Noise Modeling via Machine Learning

### v1

@@ Abstract
Maximizing the computational utility of near-term quantum processors requires predictive noise models that inform robust, noise-aware compilation and error mitigation. Conventional models often fail to capture the complex error dynamics of real hardware or require prohibitive characterization overhead. We introduce a data-efficient, machine learning-based framework to construct accurate, parameterized noise models for superconducting quantum processors. Our approach circumvents costly characterization protocols by learning hardware-specific error parameters directly from the measurement data of existing application and benchmark circuits. The generality and robustness of the framework are demonstrated through comprehensive benchmarking across multiple quantum devices and algorithms. Crucially, we show that a model trained exclusively on small-scale circuits accurately predicts the behavior of larger validation circuits. Our data-efficient approach achieves up to a 65% improvement in model fidelity quantified by the Hellinger distance between predicted and experimental circuit output distributions, compared to standard noise models derived from device properties. This work establishes a practical paradigm for noise characterization, providing a crucial tool for developing more effective noise-aware compilation and error-mitigation strategies.
@@ I Introduction
Quantum computing offers a significant opportunity to address complex optimization problems. In the current noisy intermediate-scale quantum era [1], variational algorithms exemplified by the quantum approximate optimization algorithm (QAOA) [2] and the variational quantum eigensolver (VQE) [3] have been explored for diverse applications, including portfolio optimization [4], unit commitment in power systems [5], and quantum chemistry [6]. Superconducting circuits [7, 8] have emerged as a leading platform because they benefit from established microchip manufacturing techniques and offer fast operating times. However, the practical utility of these devices is constrained by noise mechanisms and hardware limitations that hinder algorithm’s performance. To enhance the reliability of quantum applications on such devices, noise-aware compilation has become an essential strategy. Quantum compilation translates quantum algorithms into sequences of physical operations that can be executed on a target backend. This process ensures that algorithms satisfy the hardware’s intrinsic constraints and optimizes the circuit to minimize the impact of errors. Specifically, the compilation needs to adapt the algorithm to the device’s limited qubit connectivity and small set of supported gates by inserting SWAP gates and decomposing operations in algorithms into the device’s native gate set. This is enabled by compilers supported by, e.g., Qiskit [9], TKET [10] , and Munich quantum toolkit [11, 12]. Beyond ensuring basic executability, noise-aware compilation aims to maximize the fidelity of the compiled circuit. To achieve this, the compiler needs an accurate model of the device’s error characteristics, which often vary over time [13]. This model guides strategic decisions, including selecting a subset of high-fidelity qubits for execution, optimizing the routing of logical qubits to minimize noise accumulation [14, 15], and choosing between different gate sequences that, while logically equivalent, may exhibit varying resilience to noise [16] and other imperfections like Trotter errors [17, 18, 19].
@@
An accurate and predictive noise model is essential for achieving effective noise-aware compilation. However, constructing error models for real hardware presents a significant trade-off between model fidelity and experimental cost. Traditional error models, such as those offered by hardware vendors, typically assume independent Pauli noise channels that fail to capture the intricate dynamics inherent to real hardware. In practice, superconducting quantum devices exhibit spatially and temporally correlated errors [20, 21, 22], non-Markovian memory effects [23, 24, 25], and time-varying decoherence parameters [26]. Conversely, while comprehensive characterization techniques like quantum process tomography [27] can capture these intricate effects, they are resource-intensive and impractical to scale for routine calibration of larger quantum processors. Various intermediate strategies have been explored to bridge this gap. For instance, some approaches use optimization techniques such as genetic algorithms to build models for specific error types [28], while others propose scalable Markovian models that often require custom hardware tuning to be accurate [29]. Further studies have shown that noise models should be tailored to specific applications, as no single model captures all noise dynamics [30]. Even integrated approaches that combine Markovian and non-Markovian effects to improve accuracy still tend to rely on resource-intensive characterizations [31]. These studies highlight the persistent trade-off between model fidelity and experimental overhead, emphasising the necessity for a more efficient approach.
@@
To address these limitations, we introduce a practical and data-efficient alternative to resource-intensive characterization protocols. Our primary contribution is the development of a novel, circuit-size-independent parameterized noise model that achieves high fidelity by leveraging readily available experimental data, such as routine calibration measurements and standard benchmarking results, eliminating the need for dedicated quantum tomography. This approach dramatically reduces the experimental overhead typically associated with building high-fidelity models. As shown in Fig. 1, our data-efficient framework refines an initial parameterized noise model $\mathcal{N}(\bm{\theta})$ through an iterative, machine learning-driven optimization. The objective is to find optimal parameters $\bm{\theta^{*}}$ by minimizing differences between simulated and experimental output distributions, which can be quantified by the Hellinger distance [32]. A crucial feature of this approach is its ability to repurpose training data from any existing circuit execution, such as application benchmarks or algorithm runs. This eliminates the overhead of dedicated characterization protocols and allows the noise model to be tailored specifically to the context of the executed algorithms, ensuring higher fidelity for practical applications. Through extensive benchmarking across diverse quantum circuits and IBM quantum processors, we demonstrate that our optimized noise models significantly outperform the default hardware noise model.
Our framework’s predictive power and scalability are demonstrated through its performance on out-of-distribution validation circuits. Trained solely on small-scale benchmark circuits (4–6 qubits), the model accurately predicts the behavior of larger and more complex circuits (7–9 qubits). This predictive capability delivers a substantial improvement in model fidelity, reducing the mean Hellinger distance to experimental results by 50% and achieving a peak reduction of 65%. This level of predictive accuracy on larger circuits represents a significant advancement over the performance demonstrated in foundational studies on smaller systems [28]. This capability highlights the ability to learn a robust, predictive model from a small, experimentally inexpensive training set, which is critical for any scalable noise characterization protocol. The demonstrated efficacy of our framework establishes a new paradigm for noise characterization where accuracy is decoupled from high experimental overhead. Practitioners can leverage their existing experimental datasets to continuously refine noise models, enhancing the development of noise-aware compilation and error mitigation protocols [33, 34, 35] without incurring additional experimental costs.
@@
Figure 1: Flowchart of the data-efficient quantum noise modeling framework. A parameterized noise model $\mathcal{N}(\bm{\theta})$ is initially constructed based on calibration data provided by hardware vendors. The device-specific noise model $\mathcal{N}(\bm{\theta^{*}})$ is then identified by executing benchmarking circuits on real quantum hardware and aligning experimental results with noisy simulations. This iterative optimization, performed on a classical computer, leverages machine learning (ML) to update $\bm{\theta}$. This approach is highly flexible, designed to ingest outcome distributions from diverse preexisting experimental data, including application-specific circuits, avoiding the need for dedicated characterization.
@@
The paper is structured as follows. Section II presents the theoretical framework and methodology, establishing the baseline provided by device-calibrated noise models and introducing our parameterized noise model. It also details the ML technique employed for parameter optimization and concludes with a critical analysis of the framework’s scalability. Section III demonstrates a thorough experimental validation of our approach, benchmarking the model’s predictive accuracy against data from multiple superconducting processors using a diverse set of applications, including the QAOA [2], VQE [3], and random circuits. This validation includes an in-depth single-device characterization, a cross-platform comparative analysis to demonstrate hardware generality, and a study of training efficiency. Finally, Sec. IV concludes the paper.
@@ II Theoretical Framework and Methodology
This section details our noise modeling methodology. We first introduce a baseline noise model for comparison. We then introduce our circuit-size-independent parameterized noise model, followed by optimization strategy. Finally, we discuss the scalability of our approach.
@@ II.1 Default noise model
Most providers of superconducting quantum processing units (QPUs) offer noise models, which are often derived directly from the quantum device’s calibration data. A common example is the standard model available in Qiskit [9] for IBM devices, which can be constructed using the provided backend configurations and properties. This model isolates errors into two primary categories: gate errors and readout errors. Gate errors are modeled as a composed quantum channel $\mathcal{E}_{\text{gate}}$ applied after each ideal unitary operation. This channel is formed by the composition of a depolarizing channel $\mathcal{E}_{\text{depol}}$, followed by a thermal relaxation channel $\mathcal{E}_{\text{thermal}}$:
@@
Although this baseline model can be easily obtained, it excludes cross-talk and other correlated effects. Moreover, it captures only noise mechanisms explicitly characterized during routine device calibration procedures, creating a static model that cannot account for the subsequent temporal evolution of the physical device. To better capture the real noise behavior of quantum algorithms executing on QPUs, we develop a tractable, data-efficient error model, which we detail in the following section.
@@ II.2 Parameterized noise model
The model is designed to be parameterized by a vector of learnable coefficients $\bm{\theta}$, which are chosen using ML techniques to minimize the discrepancy between simulation and experimental data. An error channel $\mathcal{E}_{g}$ for any gate operation is modeled as the composition of several distinct error channels, each targeting a specific physical error mechanism.
@@ II.2.1 Single-qubit gate errors
The error channel $\mathcal{E}_{\text{1Q}}$ for any single-qubit gate (e.g., $X$, SXSX) is modeled as a composition of four channels:
$\mathcal{E}_{\text{1Q}}=\mathcal{E}_{\text{ph}}\circ\mathcal{E}_{\text{amp}}\circ\mathcal{E}_{\text{dep}}\circ\mathcal{E}_{\text{coh}}.$ (2)
The coherent error channel $\mathcal{E}_{\text{coh}}$ models systematic over- or under-rotations due to control pulse distortions. It is represented by a unitary operator $U_{\text{coh}}=\exp(-iH_{\text{coh}}t_{g})$, where $t_{g}$ is gate duration and the error Hamiltonian $H_{\text{coh}}$ is a linear combination of Pauli operators with learnable amplitudes $\theta_{x,y,z}$:
@@
The channel acts as $\mathcal{E}_{\text{coh}}(\rho)=U_{\text{coh}}~\rho~U_{\text{coh}}^{\dagger}$. For the identity gate, this coherent error is omitted.
The depolarizing error channel $\mathcal{E}_{\text{dep}}$ models a generic, incoherent disruption of the quantum state. The single-qubit depolarizing channel is given by
$\mathcal{E}_{\text{dep}}(\rho)=(1-\lambda_{\text{dep}})\rho+\frac{\lambda_{\text{dep}}}{3}(X\rho X+Y\rho Y+Z\rho Z).$ (4)
The depolarization probability $\lambda_{\text{dep}}$ is derived from the hardware-reported gate error $e_{g}$ and gate duration $t_{g}$ via a duration-aware, saturating function
@@
Here, $k_{\text{dep}}$ and $b_{\text{dep}}$ are learnable parameters that scale and offset the error rate, and $t_{\text{char}}$ is a constant characteristic gate time for the device. This form captures the physical reality that error probability saturates for long-duration operations.
The amplitude damping channel $\mathcal{E}_{\text{amp}}$ models energy relaxation. Its probability $\lambda_{\text{amp}}$ is based on the physical relaxation process for a gate of duration $t_{g}$, with an added learnable offset $b_{\text{amp}}$
@@
To capture non-Markovian dephasing, potentially arising from low-frequency noise sources like two-level-system fluctuators, we model phase damping channel $\mathcal{E}_{\text{ph}}$ using a stretched exponential decay. The probability of a phase-flip error, $\lambda_{\phi}$, is given by
@@ II.2.2 Two-qubit gate errors
The error channel for a two-qubit entangling gate, $\mathcal{E}_{\text{2Q}}$, follows a similar compositional structure but includes terms for crosstalk and correlated errors:
$\mathcal{E}_{\text{2Q}}=\mathcal{E}_{{zz}}\circ(\mathcal{E}_{\text{ph}}\otimes\mathcal{E}_{\text{ph}})\circ(\mathcal{E}_{\text{amp}}\otimes\mathcal{E}_{\text{amp}})\circ\mathcal{E}_{\text{dep}}^{(2)}\circ\mathcal{E}_{\text{coh}}^{(2)}.$ (8)
Two-qubit coherent error $\mathcal{E}_{\text{coh}}^{(2)}$ models coherent crosstalk, such as residual $ZZ$ coupling or control crosstalk. The error Hamiltonian includes two-qubit Pauli terms
@@
where $\theta_{ix,zx,zz}$ are learnable parameters. Two-qubit depolarizing channel $\mathcal{E}_{\text{dep}}^{(2)}$ introduces two learnable parameters $k_{\text{dep,2Q}}$ and $b_{\text{dep,2Q}}$, with a constant characteristic gate time $t_{\text{char,2Q}}$, forming the duration-aware, saturating probability $\lambda_{\text{dep}}^{(2)}$ analogous to the single-qubit case in Eq. (5). The single-qubit damping and dephasing channels ($\mathcal{E}_{\text{amp}},\mathcal{E}_{\text{ph}}$) are applied independently to each qubit involved in the gate, using parameters specific to the two-qubit context, including $b_{\text{amp,2Q}}$ $b_{\text{amp,2Q}}$, $\beta_{\text{2Q}}$, and $b_{\phi\text{,2Q}}$. Note that the offset parameter $b_{\phi\text{,2Q}}$ is additionally introduced for two-qubit gates, compared to single qubit gates Eq. (7), to better characterize the more complex error mechanisms of two-qubit gates.
The correlated $ZZ$ dephasing channel $\mathcal{E}_{{zz}}$ introduces a correlated dephasing error, a common mechanism in superconducting qubits arising from spectator qubit interactions or flux noise. It is modeled as a Pauli channel that applies a $Z\otimes Z$ operator with probability $\lambda_{zz}$
@@
The parameter $\lambda_{zz}$ is also modeled as a saturating, duration-dependent function parameterized by a learnable rate $k_{zz}$
$\lambda_{zz}(t_{g})=1-e^{-k_{zz}t_{g}/t_{\text{char,2Q}}}.$ (11)

### v2

@@ Abstract
Maximizing the computational utility of near-term quantum processors requires predictive noise models that inform robust, noise-aware compilation and error mitigation. Conventional models often fail to capture the complex error dynamics of real hardware or require prohibitive characterization overhead. We introduce a data-efficient framework that first constructs a physically motivated, parameterized noise model, and subsequently employs machine learning-driven Bayesian optimization to identify its parameters. Our approach circumvents costly characterization protocols by estimating algorithm- and hardware-specific error parameters directly from readily available experimental data derived from existing application and benchmark circuit executions. The generality and robustness of the framework are demonstrated across diverse algorithms and superconducting devices, yielding high-fidelity predictions by estimating an independent parameter set tailored to each specific algorithm-hardware context. Crucially, we show that a model calibrated exclusively on small-scale circuits accurately predicts the behavior of larger validation circuits. Our data-efficient approach achieves up to a 65% improvement in model fidelity quantified by the Hellinger distance between predicted and experimental circuit output distributions, compared to standard noise models derived from device properties. This work establishes a practical paradigm for application-aware noise characterization, enabling compilation and error-mitigation strategies tailored to the specific interplay between quantum algorithms and device-specific noise dynamics.
@@ I Introduction
Quantum computing offers a significant opportunity to address complex optimization problems. In the current noisy intermediate-scale quantum era [preskill2018quantum], variational algorithms exemplified by the quantum approximate optimization algorithm (QAOA) [farhi2014quantum] and the variational quantum eigensolver (VQE) [peruzzo2014variational] have been explored for diverse applications, including portfolio optimization [brandhofer2023benchmarking], unit commitment in power systems [koretsky2021adapting], and quantum chemistry [kandala2017hardware]. Superconducting circuits [clarke2008superconducting, krantz2019quantum] have emerged as a leading platform because they benefit from established microchip manufacturing techniques and offer fast operating times. However, the practical utility of these devices is constrained by noise mechanisms and hardware limitations that hinder algorithm’s performance. To enhance the reliability of quantum applications on such devices, noise-aware compilation has become an essential strategy. Quantum compilation translates quantum algorithms into sequences of physical operations that can be executed on a target backend. This process ensures that algorithms satisfy the hardware’s intrinsic constraints and optimizes the circuit to minimize the impact of errors. Specifically, the compilation needs to adapt the algorithm to the device’s limited qubit connectivity and small set of supported gates, also referred to as basis gates, by inserting SWAP gates and decomposing operations in algorithms into the device’s basis gate set. This is enabled by compilers supported by, e.g., Qiskit [qiskit], TKET [sivarajah2020tket], and Munich quantum toolkit [burgholzer2025munich, burgholzer2025MQTCore]. Beyond ensuring basic executability, noise-aware compilation aims to maximize the fidelity of the compiled circuit. To achieve this, the compiler needs an accurate model of the device’s error characteristics, which often vary over time [ji2022calibration]. This model guides strategic decisions, including selecting a subset of high-fidelity qubits for execution, optimizing the routing of logical qubits to minimize noise accumulation [quetschlich2025mqt, ji2025algorithm, montanez2025optimizing], and choosing between different gate sequences that, while logically equivalent, may exhibit varying resilience to noise [ji2023improving] and other imperfections like Trotter errors [faehrmann2022randomizing, childs2021theory, childs2019faster].
@@
An accurate and predictive noise model is essential for achieving effective noise-aware compilation. However, constructing error models for real hardware presents a significant trade-off between model fidelity and experimental cost. Traditional error models, such as those offered by hardware vendors, typically assume independent Pauli noise channels that fail to capture the intricate dynamics inherent to real hardware. In practice, superconducting quantum devices exhibit spatially and temporally correlated errors [maciejewski2021modeling, harper2020efficient, gambetta2012characterization], non-Markovian memory effects [agarwal2024modelling, zhang2022predicting, white2020demonstration], and time-varying decoherence parameters [etxezarreta2021time]. Conversely, while comprehensive characterization techniques like quantum process tomography [poyatos1997complete] can capture these intricate effects, they are resource-intensive and impractical to scale for routine calibration of larger quantum processors. Various intermediate strategies have been explored to bridge this gap. For instance, some approaches use optimization techniques such as genetic algorithms to build models for specific error types [georgopoulos2021modeling], while others propose scalable Markovian models that often require custom hardware tuning to be accurate [brand2024markovian]. Further studies have shown that noise models should be tailored to specific applications, as no single model captures all noise dynamics [dahlhauser2021modeling]. Even integrated approaches that combine Markovian and non-Markovian effects to improve accuracy still tend to rely on resource-intensive characterizations [oda2023noise]. These studies highlight the persistent trade-off between model fidelity and experimental overhead, emphasising the necessity for a more efficient approach.
@@
To address these limitations, we introduce a practical and data-efficient framework that circumvents resource-intensive characterization protocols by leveraging readily available experimental data, such as application benchmarks. Our core contribution is a circuit-size-independent parameterized noise model $\mathcal{N}(\bm{\theta})$, whose parameters are determined via iterative, machine learning-driven optimization, as depicted in Fig. 1. By minimizing the Hellinger distance [hellinger1909neue] between simulated and experimental output distributions, we estimate optimal parameters $\bm{\theta}^{*}$ directly from existing circuit executions. Crucially, this framework functions as an application-aware noise characterization. Rather than positing a single universal error profile, we estimate independent parameter sets tailored to specific algorithm-device contexts (e.g., QAOA, VQE, or random circuits on a target hardware). This approach captures the interplay between circuit structure and device-intrinsic error dynamics, yielding optimized models that significantly outperform default hardware characterizations across diverse IBM quantum processors.
We demonstrate scalability by training on 4–6 qubit benchmarks and accurately predicting out-of-distribution behavior in 7–9 qubit circuits. This approach achieves a substantial model accuracy improvement, reducing the mean Hellinger distance to experimental results by 50% (65% peak reduction). Such predictive accuracy on larger-scale systems advances beyond foundational studies restricted to small-scale demonstrations [georgopoulos2021modeling], establishing that high-fidelity, device-specific noise models can be learned from compact datasets. This enables practitioners to continuously refine noise models using existing data, advancing noise-aware compilation and error mitigation [cai2023quantum, kandala2019error, ji2024synergistic] without additional experimental cost.
@@
Figure 1: Flowchart of the data-efficient quantum noise modeling framework. A parameterized noise model $\mathcal{N}(\bm{\theta})$ is initially constructed based on calibration data provided by hardware vendors. The device-specific noise model $\mathcal{N}(\bm{\theta^{*}})$ is then identified by executing benchmarking circuits on real quantum hardware and aligning experimental results with noisy simulations. This iterative optimization, performed on a classical computer, leverages machine learning (ML)-based Bayesian optimization to update $\bm{\theta}$. This approach is highly flexible, designed to ingest outcome distributions from diverse preexisting experimental data, including application-specific circuits, avoiding the need for dedicated characterization.
@@
The paper is structured as follows. Section II presents the theoretical framework and methodology, establishing the baseline provided by device-calibrated noise models and introducing our parameterized noise model. It also details the machine learning (ML)-based Bayesian optimization (BO) technique employed for parameter determination and concludes with a critical analysis of the framework’s scalability. Section III demonstrates a thorough experimental validation of our approach, benchmarking the model’s predictive accuracy against data from multiple superconducting processors using a diverse set of applications, including the QAOA [farhi2014quantum], VQE [peruzzo2014variational], and random circuits. This validation includes an in-depth single-device characterization, a cross-platform comparative analysis to demonstrate hardware generality, and a study of training efficiency. Finally, Sec. IV concludes the paper.
@@ II Theoretical Framework and Methodology
This section details our noise modeling methodology. We first introduce a baseline noise model for comparison. We then introduce our circuit-size-independent parameterized noise model, followed by parameter determination strategy. Finally, we discuss the scalability of our approach.
@@ II.1 Default noise model
Most providers of superconducting quantum processing units (QPUs) offer noise models, which are often derived directly from the quantum device’s calibration data. These data, provided and periodically updated by the hardware vendors, quantify the intrinsic properties of the qubits, specifically their relaxation ($T_{1}$) and dephasing ($T_{2}$) times, as well as the error rates and execution durations of the device’s supported basis gates. A common example is the standard model available in Qiskit [qiskit] for IBM devices, which can be constructed using the provided backend configurations and properties. This model isolates errors into two primary categories: gate errors and readout errors. Gate errors are modeled as a composed quantum channel $\mathcal{E}_{\text{gate}}$ applied after each ideal unitary operation. This channel is formed by the composition of a depolarizing channel $\mathcal{E}_{\text{depol}}$, followed by a thermal relaxation channel $\mathcal{E}_{\text{thermal}}$:
@@
Although this baseline model can be easily obtained, it excludes cross-talk and other correlated effects. Moreover, it captures only noise mechanisms explicitly characterized during routine device calibration procedures, creating a static model that cannot account for the subsequent temporal evolution of the physical device. To better capture the real noise behavior of quantum algorithms executing on QPUs, we develop a tractable, data-efficient error model, which we detail below.
@@ II.2 Parameterized noise model
Our model is designed to be parameterized by a vector of learnable coefficients $\bm{\theta}$, which are then chosen using ML-based optimization strategy to minimize the discrepancy between simulation and experimental data. An error channel $\mathcal{E}_{g}$ for any gate operation is modeled as the composition of several distinct error channels, each targeting a specific physical error mechanism.
@@ II.2.1 Single-qubit gate errors
For a single-qubit gate $G_{q}$ (e.g., $X$, SXSX) in the bases gate set with execution duration $t_{g}$ and error rate $e_{g}$, the error channel $\mathcal{E}_{\text{1Q}}$ is modeled as a composition of four channels:
$\mathcal{E}_{\text{1Q}}=\mathcal{E}_{\text{ph}}\circ\mathcal{E}_{\text{amp}}\circ\mathcal{E}_{\text{dep}}\circ\mathcal{E}_{\text{coh}},$ (2)
where coherent errors are applied first, followed sequentially by depolarizing, amplitude damping, and phase damping channels. The coherent error channel $\mathcal{E}_{\text{coh}}$ models systematic over- or under-rotations due to control pulse distortions. This channel acts on the density matrix $\rho$ via $\mathcal{E}_{\text{coh}}(\rho)=U_{\text{coh}}~\rho~U_{\text{coh}}^{\dagger}$, defined by the unitary operator $U_{\text{coh}}=\exp(-iH_{\text{coh}}t_{g})$. The corresponding error Hamiltonian $H_{\text{coh}}$ is a linear combination of Pauli operators with learnable amplitudes $\theta_{x,y,z}$:
@@
For the identity gate, this coherent error is omitted.
The depolarizing error channel $\mathcal{E}_{\text{dep}}$ models a generic, incoherent disruption of the quantum state. For an $n$-qubit system, the depolarizing channel is given by
$\mathcal{E}_{\text{dep}}^{(n)}(\rho)=(1-\lambda_{\text{dep}}^{(n)})\,\rho+\lambda_{\text{dep}}^{(n)}\,\frac{I}{2^{n}},$ (4)
where $\lambda_{\text{dep}}^{(n)}$ is the depolarizing noise strength. For single-qubit, the noise strength $\lambda_{\text{dep}}$ is defined as a duration-aware, saturating function
@@
Here, $k_{\text{dep}}$ and $b_{\text{dep}}$ are learnable parameters that scale and offset the error rate, and $t_{\text{char}}$ is a constant characteristic gate time for the device defined as the median of all single qubit gate durations. This form captures the physical reality that error probability saturates for long-duration operations.
The amplitude damping channel $\mathcal{E}_{\text{amp}}$ models energy relaxation. Its action on the density matrix $\rho$ is defined by the Kraus decomposition
$\mathcal{E}_{\text{amp}}(\rho)=\sum_{k=0}^{1}K_{k}^{\text{amp}}\rho(K_{k}^{\text{amp}})^{\dagger},$ (6)
with Kraus operators
$K_{0}^{\text{amp}}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda_{\text{amp}}}\end{pmatrix},\quad K_{1}^{\text{amp}}=\begin{pmatrix}0&\sqrt{\lambda_{\text{amp}}}\\ 0&0\end{pmatrix}$. (7)
The relaxation probability $\lambda_{\text{amp}}$ is governed by $T_{1}$ and $t_{g}$, with an added learnable offset $b_{\text{amp}}$:
@@
To capture non-Markovian dephasing, potentially arising from low-frequency noise sources like two-level-system fluctuators, we model phase damping channel $\mathcal{E}_{\text{ph}}$ using a stretched exponential decay. Similarly, the corresponding Kraus operators are
$K_{0}^{\text{ph}}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\lambda_{\phi}}\end{pmatrix},\quad K_{1}^{\text{ph}}=\begin{pmatrix}0&0\\ 0&\sqrt{\lambda_{\phi}}\end{pmatrix}$. (9)
The dephasing parameter $\lambda_{\phi}$ is given by
@@ II.2.2 Two-qubit gate errors
For a two-qubit basis gate $G_{2Q}$ acting on the qubit pair $(q_{0},q_{1})$, with gate duration $t_{2Q}$ and error rate $e_{2Q}$, the corresponding error channel $\mathcal{E}_{\text{2Q}}$ follows a similar compositional structure but includes terms for capturing crosstalk and correlated errors:
$\mathcal{E}_{\text{2Q}}=\mathcal{E}_{{zz}}\circ(\mathcal{E}_{\text{ph}}\otimes\mathcal{E}_{\text{ph}})\circ(\mathcal{E}_{\text{amp}}\otimes\mathcal{E}_{\text{amp}})\circ\mathcal{E}_{\text{dep}}^{(2)}\circ\mathcal{E}_{\text{coh}}^{(2)},$ (11)
composing the two-qubit coherent, two-qubit depolarizing, local amplitude and phase damping, and correlated ZZ dephasing error channels. Two-qubit coherent error $\mathcal{E}_{\text{coh}}^{(2)}$ models coherent crosstalk, such as residual $ZZ$ coupling or control crosstalk. This error channel acts as $\mathcal{E}_{\text{coh}}^{(2)}(\rho)=U_{\text{coh}}^{(2)}\rho(U_{\text{coh}}^{(2)})^{\dagger}$ with $U_{\text{coh}}^{(2)}=\exp(-iH_{\text{coh}}^{(2)}t_{2Q})$. The error Hamiltonian includes two-qubit Pauli terms:
@@
where $\theta_{ix,zx,zz}$ are learnable parameters.
Analogous to the single-qubit case in Eq. (5), the two-qubit depolarizing channel $\mathcal{E}_{\text{dep}}^{(2)}$ is characterized by a duration-aware, saturating probability $\lambda_{\text{dep}}^{(2)}$, given by
$\lambda_{\text{dep}}^{(2)}(e_{2Q},t_{2Q})=1-e^{-(k_{\text{dep,2Q}}e_{g}+b_{\text{dep,2Q}})t_{2Q}/t_{\text{char,2Q}}},$ (13)
where $k_{\text{dep,2Q}}$ and $b_{\text{dep,2Q}}$ are learnable parameters. The characteristic timescale $t_{\text{char,2Q}}$ is fixed as the median duration of the two-qubit gates.
The single-qubit amplitude damping and phase dephasing channels $\mathcal{E}_{\text{amp}}$ and $\mathcal{E}_{\text{ph}}$ are applied independently to the qubit pair ($q_{0}$, $q_{1}$), using Kraus operators defined in Eqs. (7) and (9), i.e.,
$\mathcal{E}_{\text{amp}}\otimes\mathcal{E}_{\text{amp}}(\rho)=\sum_{k,l=0}^{1}(K_{k}^{\text{amp}}\otimes K_{l}^{\text{amp}})\rho(K_{k}^{\text{amp}}\otimes K_{l}^{\text{amp}})^{\dagger},$ (14)
and similarly for $\mathcal{E}_{\text{ph}}\otimes\mathcal{E}_{\text{ph}}$. Their respective error strengths $\lambda_{\text{amp},2Q}$ and $\lambda_{\phi,2Q}$ are defined as
$\lambda_{\text{amp},2Q}(t_{2Q},T_{1})$ $=1-e^{-t_{2Q}/T_{1}}+b_{\text{amp},2Q},$ (15)
$\lambda_{\phi,2Q}(t_{2Q},T_{\phi})$ $=1-e^{-2(t_{2Q}/T_{\phi})^{\beta_{2Q}}}+b_{\phi,2Q}.$ (16)
These interactions are quantified by the learnable parameters $b_{\text{amp,2Q}}$, $\beta_{\text{2Q}}$, and $b_{\phi\text{,2Q}}$. Here, the offset parameter $b_{\phi\text{,2Q}}$ is additionally introduced for two-qubit gates, compared to single qubit gates (Eq. (10)), to better characterize the more complex error mechanisms of two-qubit gates. We note that $\lambda_{\text{amp},2Q}$ and $\lambda_{\phi,2Q}$ are evaluated independently for $q_{0}$ and $q_{1}$ based on their respective coherence times ($T_{1},T_{\phi}$), yielding asymmetric local noise channels despite the common gate duration $t_{2Q}$.
The correlated $ZZ$ dephasing channel $\mathcal{E}_{{zz}}$ introduces a correlated dephasing error, a common mechanism in superconducting qubits arising from spectator qubit interactions or flux noise. It is modeled as a Pauli channel that applies a $Z\otimes Z$ operator with probability $\lambda_{zz}$:
@@
The parameter $\lambda_{zz}$ is also modeled as a saturating, duration-dependent function parameterized by a learnable rate $k_{zz}$:
$\lambda_{zz}(t_{2Q})=1-e^{-k_{zz}t_{2Q}/t_{\text{char,2Q}}}.$ (18)